OpenStudy (anonymous):

Gerald needs to purchase baseballs and bats for his youth baseball league. He’ll need at least 25 pieces of equipment, but he only has $615 to spend. Baseballs cost $20 each and bats cost $35 each. Which is a combination of baseballs and bats that Gerald can afford to buy? 16 baseballs and 9 bats 16 baseballs and 8 bats 18 baseballs and 7 bats 18 baseballs and 8 bats

OpenStudy (anonymous):

1. 16(20) + 9(35) = 320+315 = 635 No, its too much 2. 16(20) + 8(35) = 320+280 = 600 Yes, but is there one that could get more? 3. 18(20) + 7(35) = 360+ 245 = 605 Yes, you could get two more and one less 4. 18(20) + 8(35) = 360+280 = 640 No, its too much -You may want to go back and check my math and make sure its all correct but I believe that the answer is 18 baseballs and 7 bats-

OpenStudy (whpalmer4):

The problem with answer #2 is that it doesn't actually get you the 25 pieces of equipment called for in the problem statement. The problem doesn't ask that we find the answer choice that gives the most possible pieces. #2 can be immediately stricken from consideration because 16+8 < 25.